# Tracking the Feature Dynamics in LLM Training: A Mechanistic Study

### v2

@@ Appendix A Detailed Derivation of the Training-Step Continuity Theorem
Appendix B Dead Features and Ultra-Low Activation Features
@@
  • Ultra-Low Activation Features: Features with extremely low activation densities or values are also excluded. While not strictly inactive, these features exhibit negligible activations that render them semantically meaningless. This filtering is consistent with prior observations in (Bricken et al., 2023), which identify such low-activation features as non-contributive.
@@
Appendix C Weak Concept-Level Features and a Potentially Finer-Grained Concept Spectrum
@@ 5. Shelton then utilized a technique whereby …
As discussed in the main paper, there may exist a continuous spectrum and a hierarchical structure in feature levels, along with potential different geometric properties in the activation space and different training dynamics. We frame this as the Concept Spectrum Hypothesis, which suggests the feature hierarchy from lower-level token-based representations to higher-level abstract concepts. A formal quantitative characterization of this hierarchy is beyond the scope of this work and is left for future research.
Figure 7: Illustration of the Concept Spectrum Hypothesis. This figure illustrates the hierarchy from pure token-level features to higher-level concept-level features. The spectrum represents a continuous hierarchy of learned features, offering a potential refinement beyond the binary categorization discussed in our main paper.
Appendix D Polysemous Token-Level Features.
For polysemous tokens—tokens with multiple meanings—the corresponding token-level features may initially activate without capturing any semantic distinctions. During the early training phase, these features are primarily activated based on token identity alone. However, as training progresses and the model learns to incorporate semantics, these features sometimes degrade to represent only the most prominent meaning of the token. This degradation reflects the model’s learning process, where it begins to understand and refine what a token-level feature truly represents, prioritizing the most frequent or contextually significant meaning.
@@
$\Rightarrow$ 1.  North Carolina-based firm
@@
Table 3: Polysemous Token-Level Features. Words in blue denote ”resolute” or ”solid,” while green indicates ”company” or ”organization,” highlighting the model’s refinement of polysemous meanings during training.
Appendix E The Challenge of Tracking the Semantics of Initial Features
@@
  • SAE Training Property: SAE training can be viewed as selecting features from a large pool of possible features to explain the model activations (Templeton et al., 2024). Even when training SAEs twice on the same model activations and data, divergence in learned features can occur (Bricken et al., 2023). This selection process inherently introduces inconsistencies between initial and final features.
@@
  • The Impact of Possible Feature Collapse: See Discussion Appendix I.2 at Appendix I.
@@
Appendix F Implementation Details
Most experiments were conducted on a single NVIDIA A100 GPU. The implementation is built primarily upon open-source codebases (McDougall, 2024; Joseph Bloom & Chanin, 2024; Lin, 2024; Nanda & Bloom, 2022).
Models and Datasets: We use the Pythia-deduped models (Biderman et al., 2023) and Stanford CRFM Mistral (Karamcheti et al., 2021) for our experiments.
@@ Appendix F Implementation Details
Stanford CRFM Mistral (Karamcheti et al., 2021) provides an open-source replication of the GPT-2 model (Radford et al., 2019), including five GPT-2 Small and five GPT-2 Medium models trained on the OpenWebText corpus (Gokaslan et al., 2019). Each model produces 609 checkpoints, recorded at the following training steps:
@@
Stanford CRFM GPT-2 models use the OpenWebText corpus (Gokaslan et al., 2019), while Pythia models use the deduplicated version of the Pile dataset (Gao et al., 2020). The deduplicated Pile dataset ensures minimal repetition in the training data, aligning with the Pythia-deduped models.
@@
Overtraining is applied to enhance feature representations, as recommended in (Bricken et al., 2023). By leveraging the recurrent initialization scheme, which reuses pretrained weights, convergence is significantly accelerated. Specifically, only $\leq\tfrac{1}{20}$ of the initial training tokens are required for subsequent SAEs, resulting in substantial computational savings.
@@
For comparison, we also train an SAE on the final LLM checkpoint, following the commonly used approach of training SAEs on a fixed checkpoint. The results show that SAE-Track-generated SAEs exhibit similar behavior to normally trained SAEs.
@@
Appendix G Different Similarity Metrics
@@
Appendix H Experiments on Different Models
H.1 Pythia of Different Scales
@@
H.2 Stanford GPT2 of Different Scales
@@
The results are mainly consistent, except for the glitch observed in Stanford-GPT2-Small-A and the UMAP of initialization. The UMAP at initialization appears more diverged, which is related to both the initialization scheme and the model architecture. However, token-level features still exist at this stage. The glitch is further explained in Appendix I.
@@
Appendix I Possible Feature Collapse
@@
I.1 Model-Specific Phenomenon
This collapse is model-specific and may be linked to optimization settings, particularly during early training. Notably, we find that this behavior parallels observations in (Liu et al., 2022), where the cosine similarity of activations (referred to as “features” in their work but analogous to “activations” in our context) increases significantly in early training before decreasing. The observed increase in cosine similarity in our context induces the collapse, as it artificially reduces the ability of any metric to distinguish features.
I.2 Impact on SAE-Track
@@
Appendix J Research Impact and Future Directions
Here, we discuss the broader impact of our work beyond the main paper.
J.1 Broader Impact on the Field of Mechanistic Interpretability
SAE for Interpretability, Research on Training Dynamics, and Beyond:
  • Understanding what the model learns during training. Previous work (Bricken et al., 2023) has demonstrated that SAE-based interpretation not only reveals information about the data but also provides insights into the model itself, as shown by comparisons between features in randomized and trained Transformers. Our work fills the gap by systematically analyzing how features transition from those in a randomized Transformer to a fully trained Transformer, offering a mechanistic view of this process.
  • Unveiling the geometric and semantic structure of feature evolution. While SAEs trained at different checkpoints retain sparsity and can reconstruct activations in a similar manner, the extracted features—except for some token-level ones—are noisy in an untrained LLM. This is because datapoints with similar semantic meanings do not yet form distinct geometric localities in the activation space. From our geometric perspective (feature regions), feature formation becomes more intuitive, which is difficult to observe directly from conventional metrics such as reconstruction loss and sparsity. This corrects potential misconceptions that SAE structures remain unchanged during training.
  • A novel SAE training framework with potential applications to fine-tuning. The SAE-Track framework not only enables efficient tracking of features across training checkpoints but also provides a structured way to investigate changes during fine-tuning. Applying this method to fine-tuned models could reveal how learned features are modified or repurposed in response to new objectives.
J.2 Future Research Directions
Potential directions are listed below:
  • Real-Time Intervention. SAE-Track enables the synchronous training of SAEs alongside LLMs at a low computational cost, facilitating real-time supervision using SAE. Given that feature formation is a gradual and progressive process, this framework allows for the early detection and mitigation of unsafe or undesired features before they solidify. Furthermore, SAE-Track can be extended to the fine-tuning stage, providing similar capabilities for monitoring and guiding feature adaptation during model refinement. Previous studies have explored interventions or alignment using SAEs (Templeton et al., 2024; Farrell et al., 2024; Chalnev et al., 2024; Yin et al., 2024). Our approach offers a promising direction for enhancing model safety and alignment by enabling proactive feature monitoring throughout training.
  • Feature Drift and Theoretical Insights into Model Optimization. Our study reveals that feature drift persists even after features gain semantic meaning, suggesting a deeper connection between feature evolution and the model’s optimization process. Investigating this relationship could provide insights into both model training dynamics and optimization theory. Future work may explore the mechanism of how optimization methods affect drift.
  • Finer-Grained Concept Spectrum. As discussed in Appendix C, A finer-grained and quantitative analysis of this could provide deeper insights into LLM’s features.

### v3

@@ Appendix C Detailed Derivation of the Training-Step Continuity Theorem
Appendix D Feature Space Metric
Definition of Feature Space Progress Measure.
To capture a finer-grained perspective on feature formation, we extend the progress measure to the feature space itself. Given that each SAE-encoded feature captures a specific semantic representation, we define the feature space metric $M_{i}^{\text{feature}}(t)$ as:
where:
  • $\mathcal{F}_{i}^{t}$ is the set of feature vectors for the top-$k$ most activating datapoints for feature $i$ at step $t$,
  • $\mathcal{F}_{\text{random}}$ is a set of feature vectors corresponding to randomly sampled, semantically unrelated datapoints, and
  • $\overline{\text{Sim}}_{\mathcal{F}_{i}^{t}}$ and $\overline{\text{Sim}}_{\mathcal{F}_{\text{random}}}$ are the average pairwise similarities within the respective sets, defined as:
This formulation extends the datapoint-level analysis to the underlying feature vectors themselves, capturing the degree to which a feature stabilizes within its high-dimensional embedding space as training progresses.
Appendix E Dead Features and Ultra-Low Activation Features
@@
  • Ultra-Low Activation Features: Features with extremely low activation densities or values are also excluded. While not strictly inactive, these features exhibit negligible activations that render them semantically meaningless. This filtering is consistent with prior observations in [3], which identify such low-activation features as non-contributive.
@@
Appendix F Weak Concept-Level Features
@@ 5. Shelton then utilized a technique whereby …
Appendix G Polysemous Token-Level Features.
For polysemous tokens – tokens with multiple meanings – the corresponding token-level features may initially activate without capturing any semantic distinctions. During the early training phase, these features are primarily activated based on token identity alone. However, as training progresses and the model learns to incorporate semantics, these features sometimes degrade to represent only the most prominent meaning of the token. This degradation reflects the model’s learning process, where it begins to understand and refine what a token-level feature truly represents, prioritizing the most frequent or contextually significant meaning.
Table 3: Polysemous Token-Level Features. Words in blue denote “resolute” or “solid”, while green indicates “company” or “organization”, highlighting the model’s refinement of polysemous meanings during training.
@@
1.  North Carolina-based firm
@@
Appendix H The Challenge of Tracking the Semantics of Initial Features
@@
  • SAE Training Property: SAE training can be viewed as selecting features from a large pool of possible features to explain the model activations [35]. Even when training SAEs twice on the same model activations and data, divergence in learned features can occur [3]. This selection process inherently introduces inconsistencies between initial and final features.
@@
  • The Impact of Possible Feature Collapse: See Discussion Appendix L.2 at Appendix L.
@@
Appendix I Implementation Details
Most experiments were conducted on a single NVIDIA A100 GPU. The implementation is built primarily upon open-source codebases [23, 15, 20, 25].
Models and Datasets: We use the Pythia-deduped models [2] and Stanford CRFM Mistral [16] for our experiments.
@@ Appendix I Implementation Details
Stanford CRFM Mistral [16] provides an open-source replication of the GPT-2 model [30], including five GPT-2 Small and five GPT-2 Medium models trained on the OpenWebText corpus [11]. Each model produces 609 checkpoints, recorded at the following training steps:
@@
Stanford CRFM GPT-2 models use the OpenWebText corpus [11], while Pythia models use the deduplicated version of the Pile dataset [10]. The deduplicated Pile dataset ensures minimal repetition in the training data, aligning with the Pythia-deduped models.
@@
Overtraining is applied to enhance feature representations, as recommended in [3]. By leveraging the recurrent initialization scheme, which reuses pretrained weights, convergence is significantly accelerated. Specifically, only $\leq\tfrac{1}{20}$ of the initial training tokens are required for subsequent SAEs, resulting in substantial computational savings.
I.1 Comparative Study: SAE-Track vs. Conventional SAE Training
@@
For comparison, we also train an SAE on the final LLM checkpoint, following the commonly used approach of training SAEs on a fixed checkpoint. The results show that SAE-Track-generated SAEs exhibit similar behavior to normally trained SAEs, but converge significantly faster.
@@
Figure 11: Comparing converging speed between SAE-Track and normal training
I.2 Comparative Study: Reverse Tracking vs. Forward Tracking
To further validate the robustness and consistency of SAE-Track, we conducted a reverse tracking experiment. Unlike the standard forward training, where each SAE is initialized from the previous checkpoint, this approach starts from the final SAE and progressively finetunes backward through earlier checkpoints. This design aims to evaluate whether the observed feature convergence in SAE-Track is a genuine phenomenon or merely an artifact of its forward-only training strategy.
Specifically, the reverse tracking experiment addresses the concern that early-stage SAEs might prematurely consume the available capacity, limiting the ability of later stages to incorporate newly emerging, complex features. By reversing the training direction, we can test whether the observed convergence is genuinely a feature of the underlying data distribution and model architecture, rather than an unintended consequence of the forward training process.
Figures 12 and 13 present the results of this reverse tracking analysis. Despite reversing the training direction, we observe that the overall feature formation and alignment remain consistent, suggesting that the convergence observed in forward training is not solely a byproduct of incremental parameter freezing, but a more fundamental property of the model’s feature dynamics.
Figure 12: Progress Measure for reverse tracking. The top panel shows the feature space, while the bottom panel represents the activation space. Despite reversing the training sequence, the overall progression of feature formation remains consistent, indicating the stability and robustness of SAE-Track across training directions.
Figure 13: Decoder Cosine Similarity for reverse tracking. Each panel shows the alignment of decoder vectors across training checkpoints in reverse order, confirming that feature direction stability is preserved even when the training order is reversed.
Appendix J Different Similarity Metrics
@@
Appendix K Experiments on Different Models and Layers
K.1 Pythia of Other Scales
@@
K.2 Stanford GPT2 of Different Scales
@@
The results are mainly consistent, except for the glitch observed in Stanford-GPT2-Small-A and the UMAP of initialization. The UMAP at initialization appears more diverged, which is related to both the initialization scheme and the model architecture. However, token-level features still exist at this stage. The glitch is further explained in Appendix L.
@@
K.3 Experiments on Different Layers, Pythia-410m-deduped
In this section, we present experiments on different layers of the Pythia-410m-deduped model. These experiments aim to capture the feature formation and alignment dynamics across various layers, providing insights into how layer depth influences feature specialization and semantic coherence. Specifically, we include both the Progress Measure and Cosine Similarity analyses for layers 8, 12, 16, and 20. The Progress Measure captures the gradual semantic formation of features in both the activation and feature spaces, while the Cosine Similarity plots reveal the directional alignment of decoder vectors across training steps.
Figure 31: Progress Measure for Pythia-410m-deduped, layer 8. The top panel represents the activation space, while the bottom panel represents the feature space.
Figure 32: Cosine Similarity for Pythia-410m-deduped, layer 8. Each panel shows the cosine similarity between decoder vectors at different training checkpoints.
Figure 33: Feature Trajectories for Pythia-410m-deduped, layer 8.
Figure 34: Progress Measure for Pythia-410m-deduped, layer 12. The top panel represents the activation space, while the bottom panel represents the feature space.
Figure 35: Cosine Similarity for Pythia-410m-deduped, layer 12. Each panel shows the cosine similarity between decoder vectors at different training checkpoints.
Figure 36: Feature Trajectories for Pythia-410m-deduped, layer 12.
Figure 37: Progress Measure for Pythia-410m-deduped, layer 16. The top panel represents the activation space, while the bottom panel represents the feature space.
Figure 38: Cosine Similarity for Pythia-410m-deduped, layer 16. Each panel shows the cosine similarity between decoder vectors at different training checkpoints.
Figure 39: Feature Trajectories for Pythia-410m-deduped, layer 16.
Figure 40: Progress Measure for Pythia-410m-deduped, layer 20. The top panel represents the activation space, while the bottom panel represents the feature space.
Figure 41: Cosine Similarity for Pythia-410m-deduped, layer 20. Each panel shows the cosine similarity between decoder vectors at different training checkpoints.
Figure 42: Feature Trajectories for Pythia-410m-deduped, layer 20.
Appendix L Possible Feature Collapse
@@
L.1 Model-Specific Phenomenon
This collapse is model-specific and may be linked to optimization settings, particularly during early training. Notably, we find that this behavior parallels observations in [21], where the cosine similarity of activations (referred to as “features” in their work but analogous to “activations” in our context) increases significantly in early training before decreasing. The observed increase in cosine similarity in our context induces the collapse, as it artificially reduces the ability of any metric to distinguish features.
L.2 Impact on SAE-Track